OpenStudy (anonymous):

How do you write the sum of a number times 9 and 3 is equal to 6?

OpenStudy (dangerousjesse):

Sum=addition

OpenStudy (dangerousjesse):

9n+3=6

OpenStudy (anonymous):

Thank you. All that time I was writing n(9+3)=6 or 9n+3n=6